# ATM: Adversarial Tuning Multi-agent System Makes a Robust Retrieval-Augmented Generator

### v2

@@ Abstract
Large language models (LLMs) are proven to benefit a lot from retrieval-augmented generation (RAG) in alleviating hallucinations confronted with knowledge-intensive questions. RAG adopts information retrieval techniques to inject external knowledge from semantic-relevant documents as input contexts. However, due to today’s Internet being flooded with numerous noisy and fabricating content, it is inevitable that RAG systems are vulnerable to these noises and prone to respond incorrectly. To this end, we propose to optimize the retrieval-augmented Generator with a Adversarial Tuning Multi-agent system (ATM). The ATM steers the Generator to have a robust perspective of useful documents for question answering with the help of an auxiliary Attacker agent. The Generator and the Attacker are tuned adversarially for several iterations. After rounds of multi-agent iterative tuning, the Generator can eventually better discriminate useful documents amongst fabrications. The experimental results verify the effectiveness of ATM and we also observe that the Generator can achieve better performance compared to state-of-the-art baselines.
@@
Junda Zhu1††thanks: Equal contributions. Lingyong Yan211footnotemark: 1 Haibo Shi2 Dawei Yin2 Lei Sha1††thanks: Corresponding author. 1Beihang University, Beijing, China 2Baidu Inc., Beijing, China junda_zhu@outlook.com lingyongy@gmail.com shihaibo@baidu.com yindawei@acm.org shalei@buaa.edu.cn
@@ 1 Introduction
Large Language Models (LLMs) such as Llama Touvron et al. (2023a, b), Mistral Jiang et al. (2023), or GPT-4 Achiam et al. (2023) have demonstrated impressive power in massive artificial intelligence tasks, especially the question answering. However, due to the knowledge deficiency and dynamic updating, given knowledge-intensive or long-tail questions, LLMs often fail to provide concise answers, leading to rejection to answer Xu et al. (2024) or hallucination Macpherson and Platchias (2013); Zhang et al. (2023).
To alleviate this issue, Retrieval-Augmented Generation (RAG, Lewis et al., 2020) is proposed to leverage relevant external knowledge to answer open questions. Specifically, the knowledge is usually provided by relevance-based retrievers Robertson et al. (2009); Reimers and Gurevych (2019); Karpukhin et al. (2020); Gao et al. (2021), and then injected into the prompts of LLMs (termed as the Generator) to generate final answers. Most retrieval-augmented Generators refer to multiple relevant documents in practice to ensure the comprehensiveness of final answers. However, it is inevitable that retrieval systems can include some related yet useless documents or LLM-fabricating knowledge in the search results Chen et al. (2024a). As a result, the Generator in RAG systems can also suffer from incorrect and non-robust generation problems. The main reason owes to a widely observed fact that most LLMs are vulnerable to noisy or fabricating knowledge Shi et al. (2023a); Cuconasu et al. (2024); Wu et al. (2024). As illustrated in Figure 1, LLMs rely heavily on documents encountering long-tail questions, which further confirms the necessity to mitigate the impact of fabrications when generating answers.
@@
We propose a novel optimization method (termed MITO) to improve Generator’s robustness against LLM hallucination content and find it hopeful to improve the generation capacity and robustness of the Generator simultaneously.
@@ 2.1 Retrieval-Augmented Language Models
As a separate system, retriever and generator have different training objectives, giving rise to risks of indirect optimization. To this end, REPLUG Shi et al. (2023b) tunes the retriever with the output token probability without the need for direct access to generator’s parameter. RA-DIT Lin et al. (2023) introduces a co-training setting for both modules and achieves dual optimization. RAG-end2end Siriwardhana et al. (2023) proposes to dynamically re-index the document library with optimized retriever during training. Self-RAG Asai et al. (2023) and Adaptive-RAG Jeong et al. (2024) trains models to be fluent in answer generation and aware of whether to retrieve and what to retrieve, mitigating noises brought by unnecessary retrieval.
It is widely recognized that with more parameter, LLM also can act as a re-ranker better Sun et al. (2023). Recent works like MetaEOL Lei et al. (2024), GritLM Muennighoff et al. (2024), LLM2Vec BehnamGhader et al. (2024) and NV-Embed Lee et al. (2024) train or prompt LLMs to act as a sentence encoder and discover that LLMs have strong embedding capabilities. REAR Wang et al. (2024a) further plugs an extra re-ranking module to LLMs, and simultaneously improve its document ranking and answer generation capabilities.
2.2 Adversarial Learning and Robustness
Generative Adversarial Networks Goodfellow et al. (2020), widely referred to as GAN, was first proposed in image classification tasks. Its setting makes it possible that the robustness of the discriminator can be gradually enhanced with the adversarial tuning proceeding. Similar idea also works in natural language processing. Li et al. (2017) utilizes the trajectory-level reward from the discriminator distinguishing machine-generated content and conduct REINFORCE Williams (1992) to enhance the generator’s anthropomorphism.
Different from GAN-inspired methods mentioned above, we argue that, our work expands the adversarial idea to a system where both agents are generative language models. As for the RALMs, some recent works like RetRobust Yoran et al. (2023) are aimed at making LLMs robust against irrelevant documents. However, to the best of our knowledge, our proposed ATM is the first to consider the vulnerability of RAG systems to LLM fabricated knowledge and introduce and adversarial agent in the optimization of the Generator.
@@ 3.1 Attacker
With the primary goal of improving the robustness of the Generator to LLM-fabricating content, the Attacker should be able to inject fabrications into retrieved document list which can challenge the Generator successfully. In addition, some studies Liu et al. (2024) also reveal that LLMs are sensitive to positional permutations of retrieved documents. To this end, the Attacker should also be able to permute the document list to further challenge the Generator. Therefore, we devise an LLM-based Attacker that generates fabrications in two stages, namely Fabrication Generation and List Permutation.
@@
Provided queries and retrieved document lists, the Attacker generates semantically related but useless or incorrect fabrications iteratively, end up forming the attacked list containing originally retrieved documents and multiple fabricated documents.
@@ 3.2 Generator
The Generator agent (i.e. the LLM in RAG) takes the user query together with retrieved or attacked document list as the inputs, and is aimed at remaining robust to the noises and generating correct answers. To this end, a robust Generator should be capable of identifying and leveraging all relevant and useful documents for generating and ignoring those noisy knowledge, regardless of given original or fabricating document list.
@@
where denotes the language model probability Bengio et al. (2000); Radford et al. (2018) of the Generator, which is the token-level conditional probability of answer with token length as formalized in Equation 3 in detail; The is a function measuring the distance between two generation probabilities. can be further calculated by Equation. 3:
@@ 4.1 Initial Tuning
Apart from vanilla fine-tuning, we also perform three strategies to synthesize more training data based on original SFT samples: (1) answering original question with only one document, (2) answering without document, (3) ground-truth document extraction222The prompt templates can be found in Appendix A.2.. As for the Attacker, we directly prompt it without initial tuning to generate the fabrications as described in Section 3.1.
@@ 4.2 Iteratively Adversarial Optimization
where is similar to Equation 4 but uses the attacked document list as input. is the token-level Kullback–Leibler Divergence Kullback and Leibler (1951) of answer generating probabilities between given normal document list and attacked document list. is a pre-set hyper-parameter. Math derivations can be found in Appendix B, implementation details can be found in Appendix D.2.
@@ 5.1 Experimental Setup
Since most previous work Yoran et al. (2023); Asai et al. (2023); Wang et al. (2024a) uses different settings (e.g., retriever, knowledge base, number of documents included) when assessing RAG systems and there is no unified benchmark to evaluate both the generation capacity and robustness. Inspired by these studies, we conduct a novel benchmark considering both the generation capacity and robustness. The benchmark is constructed based on four main-stream RAG datasets: Natural Questions Kwiatkowski et al. (2019), TriviaQA Joshi et al. (2017), WebQuestions Berant et al. (2013) and PopQA Mallen et al. (2023).
@@
For the test set, we use the queries from the test splits of all four datasets, where PopQA is an unseen dataset for model assessment. Different from the training set, for each query, we first retrieve top-ranked documents from Wikipedia and construct some fabricated documents using powerful LLMs. Then we select top-ranked documents and fabricated documents as the final retrieved documents for each query. And we utilize Mixtral-B as the default LLM to generate fabrications.
@@
We compare our method with two state-of-the-art RALMs: REAR Wang et al. (2024a) which follows a rank-then-generate setting and Self-RAG Asai et al. (2023) which makes LLMs dynamically retrieve and generate answers. We also select a recent work–RetRobust Yoran et al. (2023), which is also aimed at improving LLMs’ robustness to irrelevant documents444We notice some contemporaneous work, i.e. RAAT Fang et al. (2024) and RobustRAG Xiang et al. (2024). Since they don’t have publicly accessible code or model weights, we didn’t reproduce their results in our experiments.. Besides, we compare our model’s performance with Llama2 model without tuning to illustrate the improvement brought by the proposed optimization.
@@ 5.2 Main Results
Table 1 shows the outperformance of ATM comparing to all baselines. Take the ATM as an example, it achieves at least Subspan EM score improvements, EM score improvements and F1 score improvements on Natural Questions dataset. We can also observe similar tendencies on the other two datasets. It verifies that through the two-stage tuning, ATM Generators can achieve better performance facing noisy retrieval documents for RAG-QA.
By comparing robustness-aware tuned models (i.e. ATM and RetRobust) with the trivially tuned RALMs (i.e. REAR and Self-RAG), we can observe considerable advancements on almost three datasets, which reveals that RALMs can benefit from robustness-aware learning to enhance its generation capacity. And ATM shows significant improvements compared with RetRobust, which serves as a strong endorsement of the effectiveness of the proposed adversarial tuning method. Besides, we find an obvious performance gap of REAR compared to its original paper encountering fabrications. The main reason can be that REAR is vulnerable to fabrications for its rank-then-generate framework. This also confirms the necessity for the robustness-aware optimization of the RALMs.
In addition, we observe that the performance increases with the ATM optimization proceeding and eventually reach the convergence after at most three iterations. It is also noteworthy that Generator without adversarial tuning (i.e., ATM) can still achieve better or comparable performance. This indicates that initial tuning is necessary to adapt Generator to the RAG-QA scenario.
@@ 5.3 Detailed Analysis
Robustness against Different Fabrication Generators
@@
Robustness against Different Fabrication Numbers
Figure 4 illustrates the robustness of different models against different fabrication numbers in ten retrieved documents (varying from to ). As is illustrated, the ATM Generator can still surpass its baselines against different fabrication numbers. Our model performs more stably (with smoother curves) when the number of fabrications increase, especially the ATM and ATM.
Performance on Unseen Dataset
To alleviate the dataset bias, we conduct experiments that the RALMs are tested with unseen dataset PopQA. The experimental results are shown in Table 3. And we can find that our methods can still surpass the baselines. Surprisingly, Self-RAG achieves the worst performance over EM and F1 scores, we induce that this may due to its self-reflection introducing extra noise of generations.
Ablation Study
We also conduct ablation to prove the necessity of different attacking types of the Attacker. As shown in Table 5, the performance drops obviously without fabrication generation (FG), which injects noisy fabrications. And the list permutation (LP) is also proven to be necessary. We induce that the positional change can increase the diversity of document list, thus enhancing the training effectiveness. We also monitored the attacking intensity of Attacker across iterations as reported in Appendix F, including a visualization analysis and the case study.
Influence of Hyper-parameter
We conduct experiments to investigate the influences of hyper-parameter settings. Specifically, we train the ATM with different in Equation 9. As is formalized, when the optimization degenerates to SFT. From Table 4 we observe that, (1) with pure SFT the optimization has relatively low performance ceilings: with a drop observed on Natural Questions; (2) lower makes the optimization more steady while a higher one brings instability. When becomes , a significant drop is witnessed at the start of optimization.
@@ 6 Conclusion
Acknowledgements
This work was supported by the National Natural Science Foundation of China under grant No. KZ37117501, No. ZG216S23E8, No. 62306024, and No. 92367204.
@@ D.1 Document Retrieval
We utilize Contriever Gao et al. (2023) 555https://huggingface.co/facebook/contriever-msmarco further fine-tuned with MS MARCO Bajaj et al. (2016) for passage retrieval. 21M Wikipedia passages dumped on Dec. 20, 2018 are adopted as our external knowledge base 666https://dl.fbaipublicfiles.com/dpr/wikipedia_split/psgs_w100.tsv.gz across datasets as listed in Section 5. In order to accelerate the vector base similarity search, we utilize faiss Johnson et al. (2019); Douze et al. (2024) and rely on GPU to accelerate the parallel search process.

### v3

@@ Abstract
Large language models (LLMs) are proven to benefit a lot from retrieval-augmented generation (RAG) in alleviating hallucinations confronted with knowledge-intensive questions. RAG adopts information retrieval techniques to inject external knowledge from semantic-relevant documents as input contexts. However, since today’s Internet is flooded with numerous noisy and fabricating content, it is inevitable that RAG systems are vulnerable to these noises and prone to respond incorrectly. To this end, we propose to optimize the retrieval-augmented Generator with an Adversarial Tuning Multi-agent system (ATM). The ATM steers the Generator to have a robust perspective of useful documents for question answering with the help of an auxiliary Attacker agent through adversarially tuning the agents for several iterations. After rounds of multi-agent iterative tuning, the Generator can eventually better discriminate useful documents amongst fabrications. The experimental results verify the effectiveness of ATM and we also observe that the Generator can achieve better performance compared to the state-of-the-art baselines. The code is available at https://github.com/chuhac/ATM-RAG.
@@
Junda Zhu1††thanks: Equal contributions. Lingyong Yan211footnotemark: 1 Haibo Shi2 Dawei Yin2 Lei Sha1††thanks: Corresponding author. 1Beihang University, Beijing, China 2Baidu Inc., Beijing, China junda_zhu@outlook.com lingyongy@gmail.com haiboshi@outlook.com yindawei@acm.org shalei@buaa.edu.cn
@@ 1 Introduction
Large Language Models (LLMs) such as Llama Touvron et al. (2023a, b); Dubey et al. (2024), Mistral Jiang et al. (2023), or GPT-4 Achiam et al. (2023) have demonstrated impressive power in massive artificial intelligence tasks, especially the question answering. However, due to knowledge deficiency and dynamic updating, given knowledge-intensive or long-tail questions, LLMs often fail to provide concise answers, leading to rejection to answer Xu et al. (2024) or hallucination Macpherson and Platchias (2013); Zhang et al. (2023).
To alleviate this issue, Retrieval-Augmented Generation (RAG, Lewis et al., 2020) is proposed to leverage relevant external knowledge to answer open questions. Specifically, the knowledge is usually provided by relevance-based retrievers Robertson et al. (2009); Reimers and Gurevych (2019); Karpukhin et al. (2020); Gao et al. (2021), and then injected into the context of LLMs (termed as the Generator) to generate final answers. Most retrieval-augmented Generators refer to multiple relevant documents in practice to ensure the comprehensiveness of final answers. However, it is inevitable that retrieval systems can include some related yet useless documents or LLM-fabricating knowledge in the search results Chen et al. (2024a). As a result, the Generator in RAG systems can also suffer from incorrect and non-robust generation problems. The main reason owes to a widely observed fact that most LLMs are vulnerable to noisy or fabricating knowledge Shi et al. (2023); Cuconasu et al. (2024); Wu et al. (2024). As illustrated in Figure 1, LLMs rely heavily on documents encountering long-tail questions, which further confirms the necessity to mitigate the impact of fabrications when generating answers.
@@
We propose a novel optimization method (termed MITO) to improve Generator’s robustness against LLM hallucination content and find it hopeful of improving the generation capacity and robustness of the Generator simultaneously.
@@ 2.1 Retrieval-Augmented Language Models
As a separate system, the retriever and generator have different training objectives, giving rise to risks of indirect optimization. To this end, REPLUG Shi et al. (2024) tunes the retriever with the output token probability without the need for direct access to generator parameters. RA-DIT Lin et al. (2024) introduces a co-training setting for both modules and achieves dual optimization. RAG-end2end Siriwardhana et al. (2023) proposes to dynamically re-index the document library with an optimized retriever during training. Self-RAG Asai et al. (2024) and Adaptive-RAG Jeong et al. (2024) trains models to be fluent in answer generation and aware of whether to retrieve and what to retrieve, mitigating noises brought by unnecessary retrieval.
It is widely recognized LLM also can act as a re-ranker better Sun et al. (2023) with more parameters. Recent works like MetaEOL Lei et al. (2024), GritLM Muennighoff et al. (2024), LLM2Vec BehnamGhader et al. (2024), NV-Embed Lee et al. (2024) and bge-en-icl Li et al. (2024) prompt, train or construct in-context learning examples for LLMs to act as a document encoder and showcase that LLMs have a strong embedding capability. REAR Wang et al. (2024a) further plugs an extra re-ranking module to LLMs, and simultaneously improves its document ranking and answer generation capabilities.
2.2 Adversarial Learning and Robust RAG
Generative Adversarial Networks Goodfellow et al. (2020), widely referred to as GAN, was first proposed in image classification tasks. Its setting makes it possible that the robustness of the discriminator can be gradually enhanced with the adversarial tuning proceeding. This idea also works in natural language processing. Li et al. (2017) utilizes the trajectory-level reward from the discriminator distinguishing machine-generated content and conduct REINFORCE Williams (1992) to enhance the generator’s anthropomorphism.
As for the RALMs, some recent works like RetRobust Yoran et al. (2024), RAAT Fang et al. (2024) and RobustRAG Xiang et al. (2024) are aimed at making LLMs robust against irrelevant and noisy documents. However, to the best of our knowledge, our proposed ATM is the first to consider the vulnerability of RAG systems to LLM fabricated content, which is prone to produce hallucinations. We introduce an adversarial agent in the optimization of the Generator on top of adversarial ideas mentioned above, and expand GAN-inspired methods to a system where both agents are generative language models.
@@ 3.1 Attacker
With the primary goal of improving the robustness of the Generator to LLM-fabricating content, the Attacker should be able to inject fabrications into the retrieved document list which can challenge the Generator successfully. In addition, some studies Liu et al. (2024) also reveal that LLMs are sensitive to positional permutations of retrieved documents. To this end, the Attacker should also be able to permute the document list to further challenge the Generator. Therefore, we devise an LLM-based Attacker that generates fabrications in two stages, namely Fabrication Generation and List Permutation.
@@
Provided queries and retrieved document lists, the Attacker generates semantically related but useless or incorrect fabrications iteratively, and ends up forming the attacked list containing originally retrieved documents and multiple fabricated documents.
@@ 3.2 Generator
The Generator agent (i.e. the LLM in RAG) takes the user query together with retrieved or attacked document list as the inputs, and is aimed at remaining robust to the noises and generating correct answers. To this end, a robust Generator should be capable of identifying and leveraging all relevant and useful documents for generating and ignoring noisy knowledge, regardless of given original or fabricating document list.
@@
where denotes the language model probability Bengio et al. (2000); Radford et al. (2018) of the Generator, which is the token-level conditional probability of answer with token length as formalized in Equation 3 in detail; The is a function measuring the distance between two generation probabilities. can be further calculated by Eqn. 3:
@@ 4.1 Initial Tuning
Apart from vanilla fine-tuning, we also perform three strategies to synthesize more training data based on original SFT samples: (1) answering the original question with only one document, (2) answering without document, (3) ground-truth document extraction222The prompt templates can be found in Appendix A.2.. As for the Attacker, we directly prompt it without initial tuning to generate the fabrications as described in Section 3.1.
@@ 4.2 Iteratively Adversarial Optimization
where is similar to Equation 4 but uses the attacked document list as input. is the token-level Kullback–Leibler Divergence Kullback and Leibler (1951) of answer generating probabilities between the given normal document list and the attacked document list. is a pre-set hyper-parameter. Math derivations can be found in Appendix B, and implementation details can be found in Appendix D.2.
@@ 5.1 Experimental Setup
Since most previous work Yoran et al. (2024); Asai et al. (2024); Wang et al. (2024a) uses different settings (e.g., retriever, knowledge base, number of documents included) when assessing RAG systems, there is no unified benchmark to evaluate both the generation capacity and robustness. Inspired by these studies, we conduct a novel benchmark considering both the generation capacity and robustness. The benchmark is constructed based on four main-stream RAG datasets: Natural Questions Kwiatkowski et al. (2019), TriviaQA Joshi et al. (2017), WebQuestions Berant et al. (2013) and PopQA Mallen et al. (2023).
@@
For the test set, we use the queries from the test splits of all four datasets, where PopQA is an unseen dataset during training, for model assessment. Different from the training set, for each query, we first retrieve top-ranked documents from Wikipedia and construct some fabricated documents using powerful LLMs. Then we select top-ranked documents and fabricated documents as the final retrieved documents for each query. And we utilize Mixtral-B as the default LLM to generate fabrications.
@@
We compare our method with four state-of-the-art RALMs: 1) REAR Wang et al. (2024a) which follows a rank-then-generate setting; 2) Self-RAG Asai et al. (2024) which makes LLMs self-perceptively retrieve external knowledge and generate answers; 3) RetRobust Yoran et al. (2024) which is aimed at improving LLMs’ robustness to irrelevant documents; 4) RAAT Fang et al. (2024) which enhances LLMs’ performance to generate answers and discriminate noisy documents through dual-task learning on the constructed dataset.
@@ 5.2 Main Results
Table 1 shows the outperformance of ATM comparing to all baselines. Take the ATM as an example, it achieves at least Subspan EM score improvements, EM score improvements and F1 score improvements on Natural Questions dataset. We can also observe similar tendencies on the other datasets. It verifies that through the two-stage tuning, ATM Generators can achieve better performance when facing noisy retrieval documents for RAG-QA.
By comparing robustness-aware tuned models (i.e. ATM, RetRobust and RAAT) with the trivially tuned RALMs (i.e. REAR and Self-RAG), we can observe considerable advancements on almost three datasets, which reveals that RALMs can benefit from robustness-aware learning to enhance their generation capacity. And ATM usually achieves significant improvements compared with RetRobust and RAAT, which serves as a strong endorsement of the effectiveness of the proposed adversarial tuning method. Besides, we find an obvious performance gap of REAR compared to its original paper encountering fabrications. The main reason can be that REAR is vulnerable to fabrications for its rank-then-generate framework. This also confirms the necessity for the robustness-aware optimization of the RALMs.
In addition, we can also observe that the performance increases with the ATM optimization proceeding and eventually reach the convergence after at most iterations. It is also noteworthy that Generator without adversarial tuning (i.e., ATM) can still achieve better or comparable performance. This indicates that initial tuning is necessary to adapt Generator to the RAG-QA scenario.
@@ 5.3 Detailed Analysis
Robustness against different fabrication generators.
@@
Robustness against different fabrication numbers.
Figure 4 illustrates the robustness of different models against different fabrication numbers (varying from to ) in the case of a total of documents. As is illustrated, the ATM Generator has more stable performance (with smoother curves) when the number of fabrications increase, especially the ATM and ATM. When there is less fabrications, our model also performs comparably with state-of-the-art RALMs. And with the number of fabrications increasing, our model can surpass its baselines, showcasing its stability and robustness.
Performance on unseen dataset.
To alleviate the dataset bias, we conduct experiments to evaluate the RALMs on an unseen dataset–PopQA. The experimental results are shown in Table 3. And we can find that our methods can still surpass the baselines. Surprisingly, Self-RAG achieves the worst performance over EM and F1 scores, we induce that this may be due to its self-reflection introducing extra noise of generations.
Visualization of attacking intensity.
We also monitored the attacking intensity of Attacker as optimization proceeds. For the visualization analysis, we analyze the Log Loss of Generator given Attacker fabrications at different iterations. As is showcased in Figure 5, the increasing PPL of the Generator shows that Attacker is emerging increasingly stronger attacking patterns which greatly obstruct the Generator from generating right answers. In order to observe the change of attacking patterns more intuitively, we also conduct case study as reported in Appendix F.
Ablation study.
We also conduct ablation to verify the necessity of different attacking types of the Attacker. As shown in Table 5, the performance drops obviously without fabrication generation (FG), which injects noisy fabrications. The list permutation (LP) is also proven to be necessary. We induce that the positional change can increase the diversity of the document list, thus enhancing the training effectiveness.
Influence of hyper-parameters.
We conduct experiments to investigate the influences of hyper-parameter settings. Specifically, we train the ATM with different in Equation 9. As is formalized, when the optimization degenerates to SFT. From Table 4 we observe that, (1) with vanilla SFT the optimization has relatively low-performance ceilings: with a drop observed on Natural Questions; (2) lower makes the optimization more steady while a higher one brings instability. When becomes , a significant drop is witnessed at the start of optimization.
@@ 6 Conclusion
Acknowledgement
This work was supported by the National Science Fund for Excellent Young Scholars (Overseas) under grant No. KZ37117501, National Natural Science Foundation of China (No. 62306024, No. 92367204).
@@ D.1 Document Retrieval
We utilize Contriever Gao et al. (2023) 444https://huggingface.co/facebook/contriever-msmarco further fine-tuned with MS MARCO Bajaj et al. (2016) for passage retrieval. 21M Wikipedia passages dumped on Dec. 20, 2018 are adopted as our external knowledge base 555https://dl.fbaipublicfiles.com/dpr/wikipedia_split/psgs_w100.tsv.gz across datasets as listed in Section 5. In order to accelerate the vector base similarity search, we utilize faiss Johnson et al. (2019); Douze et al. (2024) and rely on GPU to accelerate the parallel search process.